OpenStudy (ooxbrittanyxoo):

The formula to find the period of orbit of a satellite around a planet is T^2=(4pi^2/GM)r^3 where r is the orbit's mean radius, M is the mass of the planet, and G is the universal gravitational constant. if you are given all the values except r, how do you rewrite the formula to solve for r? will medal

OpenStudy (ooxbrittanyxoo):

A) \[r=\sqrt[3]{GMT \div4\pi ^{2}}\] B) \[r=T \div2\pi \sqrt[3]{GM}\] C) \[r=\sqrt[3]{GMT ^{2}\div4\pi ^{2}}\] D) \[r=\sqrt[3]{GMT ^{2}\div4\pi}\]

syed98 (syedmohammed98):

http://openstudy.com/study#/updates/55b26bffe4b0ce10565f633c there ya go

OpenStudy (ooxbrittanyxoo):

thnx lots